An iron rod 4 m long and 0.5 cm2 in cross section stretches 1 mm when a mass of 225 kg is hung from its lower end. Compute Young’s modulus for the iron.


1
Expert's answer
2021-07-09T08:31:35-0400
"\frac{mg}{A}=Y\frac{\Delta L}{L}\\\frac{(225)(9.81)}{0.5\cdot10^{-4}}=Y\frac{0.003}{4}\\Y=177\ GPa"